OpenStudy (hba):

What is the distance between the two intercepts of the graph y=x^2-12x+20.

OpenStudy (hba):

@ash2326 Undefined.

OpenStudy (hba):

@ajprincess You are welcome to give hints...

OpenStudy (anonymous):

8

OpenStudy (hba):

What will i do of the answer lol haha @MarinaDL

OpenStudy (hba):

Tell me how to do it lol ?

OpenStudy (anonymous):

why lol? or does it mean thank you? x=2 and 10

OpenStudy (hba):

@MarinaDL I do not want the answer i know the answer i just want you to propose a solution if you have ?

OpenStudy (hba):

@MarinaDL Just give me hints lol

OpenStudy (anonymous):

I already did my best.

OpenStudy (hba):

You did not help at all,you are just violating the CoC- http://openstudy.com/code-of-conduct

OpenStudy (hba):

@ash2326 Pleaseeee Help :D

OpenStudy (anonymous):

when x=0 , y=0^2 -12*0+20 =20 So, y-intercept = 20 when y=0 , 0=x^2-12x+20 x = 2 or 10 So, x-intercept = 2 and 10

OpenStudy (hba):

Good going what next ?

OpenStudy (anonymous):

And u want distance between which two intercepts?

OpenStudy (hba):

x intercepts ?

OpenStudy (anonymous):

Thank you for sharing with me C-O-C. Sorry for the direct answer. What about "I will thank people for taking the time and effort to help me" from your side?

OpenStudy (hba):

You were not helping you should go through my last ques @MarinaDL . @waterineyes and @hartnn never give me the answer,they just hint me to it and i give them medals and thank them,I do not need answers or total solutions because i do not have to show my work neither i have to just get the answer. I am just revising my work,if you hint me like they do i will defiantly give you a medal and be very thankful to you.Sorry @MarinaDL if you think i came of as rude.

OpenStudy (anonymous):

@hba u want distance between two x-intercepts?

OpenStudy (hba):

Yeah :)

OpenStudy (hba):

10-2=8

OpenStudy (anonymous):

correct